# Ikeda type lift on ${\rm SO}(3,n+1)$

### v2

@@ Abstract.
By using Ikeda’s theory for a compatible family of Eisenstein series, we explicitly construct Ikeda type lifts on the special orthogonal group $G={\rm SO}(3,n+1)$ over $\mathbb{Q}$ with $n\geq 3$ which splits everywhere at finite places. Our lifts are Hecke eigen cusp forms of weight $l$ ($l\geq n+2$, even) and come from elliptic newforms with respect to ${\rm SL}_{2}(\mathbb{Z})$ which are of weight $l-\frac{n-2}{2}$ when $n$ is even and 2ln+12l-n+1 when $n$ is odd.
@@ 1. Introduction
Let $\mathbb{A}=\mathbb{A}_{\mathbb{Q}}$ be the ring of adeles of $\mathbb{Q}$. For an integer $n\geq 3$, let $A$ be a positive definite integral matrix of size n2n-2. Assume $A$ is even (namely, any diagonal entries are even integers) if $n$ is even while $A={\rm diag}(1,A^{\prime})$ if $n$ is odd where $A^{\prime}$ is a positive definite integral matrix of size n3n-3. Let $J_{2,n}={\rm antidiag}(1,1,-A,1,1)$ and $J_{3,n+1}={\rm antidiag}(1,1,1,-A,1,1,1)$ which are of size $n+3$ and $n+4$ respectively (see Section 2). Let $G={\rm SO}(3,n+1)$ be the special orthogonal group over $\mathbb{Q}$ associated to the symmetric pairing defined by $J_{3,n+1}$. Note that if $n$ is even, $G$ does not have discrete series. Assume $G$ splits everywhere at finite places. By [26, Section 2.1], it is equivalent to $\det(A)=1$ and $n\equiv 2$ mod 8 when $n$ is even, $\det(A^{\prime})=1$ and $n\equiv 3$ mod 8 when $n$ is odd. For instance, this condition is satisfied when $n=8a+2$, $A={\rm diag}(\overbrace{A_{8},\ldots,A_{8}}^{a})$ or $n=8a+3$, $A={\rm diag}(1,\overbrace{A_{8},\ldots,A_{8}}^{a})$ where $A_{8}$ is the $E_{8}$ Cartan matrix given as an element of even integral matrix in $M_{8}(\mathbb{Z})$. This agrees with the result of [8] since $\frac{n+4}{2}\equiv 3$ (mod 4) if $n$ is even, and $\frac{n+3}{2}\equiv 3$ (mod 4).
Let $P$ be the Siegel parabolic subgroup of $G$ with the Levi decomposition $P=MN$ where $M=\{{\rm diag}(t,m,t^{-1})\ |\ t\in{\rm GL}_{1},\ m\in{\rm SO}(2,n)\}\simeq{\rm GL}_{1}\times{\rm SO}(2,n)$ and $N\simeq\mathbb{G}^{n+2}_{a}$ where ${\rm SO}(2,n)$ is defined by $J_{2,n}$. Let $V^{\prime}$ be the quadratic space associated to $J_{2,n}$ with the quadratic map $q:V^{\prime}\longrightarrow\mathbb{G}_{a}$ and we identify $V^{\prime}$ with $N$. We also denote by $(\ast,\ast)_{V^{\prime}}$ the corresponding symmetric bilinear pairing.
In [23, Section 4], Pollack defined the Eisenstein series $E_{l}$ on $G(\mathbb{A})$ of weight $l$ with respect to G(^)G($\widehat{\mathbb{Z}}$), and computed unramified and archimedean part of Fourier coefficients. He can show the algebraicity of the Fourier coefficients. However, he did not compute the Siegel series explicitly which come from the bad finite places. The computation of the Siegel series is complicated and it is contained in Section 11. One corollary of our explicit computation is that we can show that Fourier coefficients of Eisenstein series have bounded denominator. (See Theorem 4.8 and 4.9.) After a long computation, we can write down the Fourier expansion of the Eisenstein series as follows:
@@
where $\mathcal{W}_{2\pi\eta}$ is Pollack’s spherical function associated to the additive character $e^{2\pi\sqrt{-1}(\eta,\ast)_{V^{\prime}}}$ on $V^{\prime}(\mathbb{R})$ (see [23, Section 3] but we use a slightly different normalization as explained later). We remark that $a_{E_{l}}(\eta)=0$ unless $\eta\in V^{\prime}(\mathbb{Z})$.
@@
where the constant $C_{l,n}$ is given explicitly in Theorem 4.8. We remark that $\widetilde{Q}_{\eta,\Phi_{p}}(X_{p})$ is also defined for any $\eta\in V^{\prime}(\mathbb{Q}_{p})$ with $q(\eta)\neq 0$. For any $g_{p}={\rm diag}(t_{p},m_{p},t^{-1}_{p})n_{p}\in P(\mathbb{Q}_{p})=M(\mathbb{Q}_{p})N(\mathbb{Q}_{p})$, we will see later that $\widetilde{Q}_{\eta,g_{p}\Phi_{p}}(X_{p})=|t_{p}|^{\frac{n+2}{2}}_{p}\widetilde{Q}_{g^{-1}_{p}\eta,\Phi_{p}}(X_{p})$. Thus, $\widetilde{Q}_{\eta,g_{p}\cdot\Phi_{p}}(X_{p})$ is also a Laurent polynomial (over $\mathbb{Q}$) satisfying $\widetilde{Q}_{\eta,g_{p}\cdot\Phi_{p}}(X_{p})=\widetilde{Q}_{\eta,g_{p}\cdot\Phi_{p}}(X^{-1}_{p})$.
@@
$F_{f}(g)=\sum_{\eta\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0}A_{f}(\eta)(g_{\mathbb{f}})\mathcal{W}_{2\pi\eta}(g_{\infty}),\ g=g_{\mathbb{f}}g_{\infty}\in G(\mathbb{A})$
is a non-zero Hecke eigen cusp form of weight $l$ with respect to G(^)G($\widehat{\mathbb{Z}}$). Let $\Pi_{F_{f}}=\otimes^{\prime}\Pi_{p}$ be the cuspidal representation associated with $F_{f}$. Further, the standard $L$-function of $\Pi_{F_{f}}$ is given by
@@
$F_{f}(g)=\sum_{\eta\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0}A_{f}(\eta)(g_{\mathbb{f}})\mathcal{W}_{2\pi\eta}(g_{\infty}),\ g=g_{\mathbb{f}}g_{\infty}\in G(\mathbb{A})$
is a non-zero Hecke eigen cusp form of weight $l$ with respect to G(^)G($\widehat{\mathbb{Z}}$). Let $\Pi_{F}=\otimes^{\prime}\Pi_{p}$ be the cuspidal representation associated with $F_{f}$. Further, the standard $L$-function of $\Pi_{F_{f}}$ is given by
$L(s,\Pi_{F_{f}})=L(s,\pi_{f})\prod_{i=0}^{n}\zeta(s+\frac{n}{2}-i).$
@@
Acknowledgments. We would like to thank Hiroaki Narita and Jim Arthur for helpful discussions. We thank KIAS in Seoul and Waseda University in Tokyo for their incredible hospitality during this research.
@@ 2.2. Another maximal parabolic subgroup and a Jacobi group
as in [15, (4.5),p.231]. Thus, we have the Jacobi group in our setting:
@@ 3. Automorphic forms on $G={\rm SO}(3,n+1)$
Recall $G={\rm SO}(V)$. The maximal compact subgroup $K$ of $G(\mathbb{R})$ is $S({\rm O}(3)\times{\rm O}(n+1))$. We have a map $S({\rm O}(3)\times{\rm O}(n+1))\rightarrow{\rm O}(3)=({\rm SU}(2)/\mu_{2})\rtimes\{\pm 1\}$. Denote by $\mathbb{V}_{l}={\rm Sym}^{2l}(\mathbb{C}^{2})$ the $(2l+1)$-dimensional representation of $K$ that factors through ${\rm O}(3)$. We fix a basis $\{[x]^{l+v}[y]^{l-v}\}_{v=-l}^{l}$ of $V_{l}$ where $x^{k}]=\frac{x^{k}}{k!}$ and $y^{k}]=\frac{y^{k}}{k!}$ for $k\geq 0$. Modular forms on $G$ of weight $l$ are $\mathbb{V}_{l}$-valued automorphic functions $\phi$ on $G(\mathbb{A})$ that satisfy:
@@
Definition 3.2.
We say that $\phi$ is a cusp form of level one if it has the Fourier expansion of the form
$\phi(g)=\sum_{\eta\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0}a_{\phi}(\eta)(g_{f})\mathcal{W}_{2\pi\eta}(g_{\infty}).$

### v3

@@ Abstract.
By using Ikeda’s theory for a compatible family of Eisenstein series, we explicitly construct Ikeda type lifts on the special orthogonal group $G={\rm SO}(3,n+1)$ over $\mathbb{Q}$ with $n\geq 3$ which splits everywhere at finite places. Our lifts are Hecke eigen cusp forms of weight $l$ ($l\geq n+2$, even) and come from elliptic newforms with respect to ${\rm SL}_{2}(\mathbb{Z})$ which are of weight $l-\frac{n-2}{2}$ when $n$ is even and 2ln+12l-n+1 when $n$ is odd. The corresponding cuspidal automorphic representations are cohomological but non-tempered at any places including the infinity place.
@@ 1. Introduction
Let $\mathbb{A}=\mathbb{A}_{\mathbb{Q}}$ be the ring of adeles of $\mathbb{Q}$. For an integer $n\geq 3$, let $A$ be a positive definite integral matrix of size n2n-2. Assume $A$ is even (namely, any diagonal entries are even integers) if $n$ is even while $A={\rm diag}(1,A^{\prime})$ if $n$ is odd where $A^{\prime}$ is a positive definite integral matrix of size n3n-3. Let $J_{2,n}={\rm antidiag}(1,1,-A,1,1)$ and $J_{3,n+1}={\rm antidiag}(1,1,1,-A,1,1,1)$ which are of size $n+3$ and $n+4$ respectively (see Section 2). Let $G={\rm SO}(3,n+1)$ be the special orthogonal group over $\mathbb{Q}$ associated to the symmetric pairing defined by $J_{3,n+1}$. Note that if $n$ is even, $G$ does not have discrete series. Assume $G$ splits everywhere at finite places. By [32, Section 2.1], it is equivalent to $\det(A)=1$ and $n\equiv 2$ mod 8 when $n$ is even, $\det(A^{\prime})=1$ and $n\equiv 3$ mod 8 when $n$ is odd. For instance, this condition is satisfied when $n=8a+2$, $A={\rm diag}(\overbrace{A_{8},\ldots,A_{8}}^{a})$ or $n=8a+3$, $A={\rm diag}(1,\overbrace{A_{8},\ldots,A_{8}}^{a})$ where $A_{8}$ is the $E_{8}$ Cartan matrix given as an element of even integral matrix in $M_{8}(\mathbb{Z})$. This agrees with the result of [12] since $\frac{n+4}{2}\equiv 3$ (mod 4) if $n$ is even, and $\frac{n+3}{2}\equiv 3$ (mod 4).
Let $P$ be the Siegel parabolic subgroup of $G$ with the Levi decomposition $P=MN$ where $M=\{{\rm diag}(t,m,t^{-1})\ |\ t\in{\rm GL}_{1},\ m\in{\rm SO}(2,n)\}\simeq{\rm GL}_{1}\times{\rm SO}(2,n)$ and $N\simeq\mathbb{G}^{n+2}_{a}$ where ${\rm SO}(2,n)$ is defined by $J_{2,n}$. Let $V^{\prime}$ be the quadratic space associated to $J_{2,n}$ with the quadratic map $q:V^{\prime}\longrightarrow\mathbb{G}_{a}$ and we identify $V^{\prime}$ with $N$. We also denote by $(\ast,\ast)_{V^{\prime}}$ the corresponding symmetric bilinear pairing. Let $V^{\prime\prime}$ be the quadratic space associated to $J_{1,n-1}={\rm antidiag}(1,-A,1)$ and embed $V^{\prime\prime}$ into $V^{\prime}$ by $\xi\mapsto(0,\xi,0)$. We write any element of $V^{\prime}(\mathbb{Q})$ as $\eta=(a,\xi,S)$ with $a,S\in\mathbb{Q}$ and $\xi\in V^{\prime\prime}(\mathbb{Q})$.
In [29, Section 4], Pollack defined the Eisenstein series $E_{l}$ on $G(\mathbb{A})$ of weight $l$ with respect to G(^)G($\widehat{\mathbb{Z}}$), and computed unramified and archimedean part of Fourier coefficients. He can show the algebraicity of the Fourier coefficients. However, he did not compute the Siegel series explicitly which come from the bad finite places. The computation of the Siegel series is complicated and it is contained in Section 11. One corollary of our explicit computation is that we can show that Fourier coefficients of Eisenstein series have bounded denominator. (See Theorem 4.8 and 4.9.) After a long computation, we can write down the Fourier expansion of the Eisenstein series as follows:
@@
where $\mathcal{W}_{2\pi\eta}$ is Pollack’s spherical function associated to the additive character $e^{2\pi\sqrt{-1}(\eta,\ast)_{V^{\prime}}}$ on $V^{\prime}(\mathbb{R})$ (see [29, Section 3] but we use a slightly different normalization as explained later). We remark that $a_{E_{l}}(\eta)=0$ unless $\eta\in V^{\prime}(\mathbb{Z})$.
@@
where the constant $C_{l,n}$ is given explicitly in Theorem 4.8. We remark that $\widetilde{Q}_{\eta,\Phi_{p}}(X_{p})$ is also defined for any $\eta\in V^{\prime}(\mathbb{Q}_{p})$ with $q(\eta)\neq 0$. For any $g_{p}={\rm diag}(t_{p},m_{p},t^{-1}_{p})n_{p}\in P(\mathbb{Q}_{p})=M(\mathbb{Q}_{p})N(\mathbb{Q}_{p})$, we will see later that $\widetilde{Q}_{\eta,g_{p}\cdot\Phi_{p}}(X_{p})=|t_{p}|^{\frac{n+2}{2}}_{p}\widetilde{Q}_{g^{-1}_{p}\eta,\Phi_{p}}(X_{p})$. Thus, $\widetilde{Q}_{\eta,g_{p}\cdot\Phi_{p}}(X_{p})$ is also a Laurent polynomial (over $\mathbb{Q}$) satisfying $\widetilde{Q}_{\eta,g_{p}\cdot\Phi_{p}}(X_{p})=\widetilde{Q}_{\eta,g_{p}\cdot\Phi_{p}}(X^{-1}_{p})$.
@@
$F_{f}(g)=\sum_{\eta=(a,\xi,S)\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0,\ aS\neq 0}A_{f}(\eta)(g_{\mathbb{f}})\mathcal{W}_{2\pi\eta}(g_{\infty}),\ g=g_{\mathbb{f}}g_{\infty}\in G(\mathbb{A})$
is a non-zero Hecke eigen cusp form of weight $l$ with respect to G(^)G($\widehat{\mathbb{Z}}$). Let $\Pi_{F_{f}}=\Pi_{\infty}\otimes\otimes^{\prime}_{p}\Pi_{p}$ be the cuspidal representation associated with $F_{f}$. Then the standard $L$-function of $\Pi_{F_{f}}$ is given by
@@
Further, as a $(\mathfrak{g},K)$-module, $\Pi_{\infty}$ is isomorphic to $\Pi^{3,n+1}_{l,0}$ in the notation of [26, Proposition 2.1] which is a cohomological, non-tempered representation.
@@
$F_{f}(g)=\sum_{\eta=(a,\xi,S)\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0,\ aS\neq 0}A_{f}(\eta)(g_{\mathbb{f}})\mathcal{W}_{2\pi\eta}(g_{\infty}),\ g=g_{\mathbb{f}}g_{\infty}\in G(\mathbb{A})$
is a non-zero Hecke eigen cusp form of weight $l$ with respect to G(^)G($\widehat{\mathbb{Z}}$). Let $\Pi_{F_{f}}=\Pi_{\infty}\otimes\otimes^{\prime}_{p}\Pi_{p}$ be the cuspidal representation associated with $F_{f}$. Then the standard $L$-function of $\Pi_{F_{f}}$ is given by
$L(s,\Pi_{F_{f}})=L(s,\pi_{f})\prod_{i=0}^{n}\zeta\Big(s+\frac{n}{2}-i\Big).$
Further, as a $(\mathfrak{g},K)$-module, $\Pi_{\infty}$ is isomorphic to $\Pi^{3,n+1}_{l,0}$ in the notation of [26, Proposition 2.1] which is a cohomological, non-tempered representation.
Remark 1.3.
The global A-parameters associated with our lifts $\Pi_{F_{f}}=\Pi_{\infty}\otimes\otimes^{\prime}_{p}\Pi_{p}$ are similar to those of the theta lifts from ${\rm SL}_{2}$ or its double covering $\widetilde{{\rm SL}_{2}}$ to ${\rm O}(3,n+1)$ (cf. [2, p. 360, Theorem 4.3] or [9, p.4789, (4.2)]). They differ by the twist of the quadratic character $\chi_{V}$ (note that $\det(V)=-1$).
On the other hand, it follows from [36, p. 712, Proposition 10.1(3),(4)] (by taking j=3,n=2j=3,\ n=2, $j_{0}\leq 2$, and $\rho_{2}=3$ therein) that the theta lift from $f$ produces a non-zero, square-integrable automorphic forms on ${\rm O}(3,n+1)$. Applying [36, p. 712, Proposition 10.1(3)] again with $j=2$ and $n=2$, together with the Rallis tower property, we see that such a lift is non-cuspidal, which is consistent with the situation described in [26]. Further, the theta lifts have a non-trivial level by the appearance of $\chi_{V}$.
In Proposition 3.2, we show that $\Pi_{\infty}\simeq\Pi^{3,n+1}_{l,0}$, and it is realized as a subrepresentation of the degenerate principal series representation ${\rm Ind}_{P(\mathbb{R})}^{G(\mathbb{R})}\,|\nu|^{l-\frac{n}{2}}$. On the other hand, the infinity component of the theta lift should be an irreducible constituent of ${\rm Ind}_{P(\mathbb{R})}^{G(\mathbb{R})}\,{\rm sgn}(\nu)|\nu|^{l-\frac{n}{2}}$. We give the details in our forthcoming paper [21].
As an immediate result with [33, p.64, THEOREM 3.3], we obtain the following corollary:
Corollary 1.4.
Keep the notation as above. Let $\Gamma:={\rm SO}(3,n+1)(\mathbb{Z})$ and $X=G(\mathbb{R})/K$ be the associated symmetric space of real dimension $3(n+1)$, where $K$ is a maximal compact subgroup of $G(\mathbb{R})$. Then $\Pi_{F_{f}}$ contributes to the cuspidal part $H^{n+1}_{{\rm cusp}}(\Gamma\backslash X,\mathcal{E}_{l})$ of the cohomology $H^{n+1}(\Gamma\backslash X,\mathcal{E}_{l})$, where $\mathcal{E}_{l}$ is the local system associated with the algebraic representation $E_{l}={\rm Sym}^{l}(\mathbb{C}^{n+4})_{{\rm harm}}$ of $G(\mathbb{R})$. Here the subscript “harm” means the harmonic polynomials in ${\rm Sym}^{l}(\mathbb{C}^{n+4})$ with respect to the variables of $\mathbb{C}^{n+4}$.
We refer to [24] and [3] for the (cuspidal) cohomology groups of the arithmetic manifolds associated to orthogonal groups. A proof of the above corollary is given in Section 8.5.
@@
Acknowledgments. We would like to thank Tamotsu Ikeda, Hiraku Atobe, Hiroaki Narita and Jim Arthur for helpful discussions, and Yi Shan for pointing out some inaccuracies. We thank KIAS in Seoul and Waseda University in Tokyo for their incredible hospitality during this research.
@@ 2.2. Another maximal parabolic subgroup and a Jacobi group
as in [19, (4.5),p.231]. Thus, we have the Jacobi group in our setting:
@@ 3. Automorphic forms on $G={\rm SO}(3,n+1)$
Recall $G={\rm SO}(V)$. The maximal compact subgroup $K$ of $G(\mathbb{R})$ is $S({\rm O}(3)\times{\rm O}(n+1))$. Let $K_{0}={\rm SO}(3)\times{\rm SO}(n+1)$. We have a map $S({\rm O}(3)\times{\rm O}(n+1))\rightarrow{\rm O}(3)=({\rm SU}(2)/\mu_{2})\rtimes\{\pm 1\}$. For each $\lambda\in\mathbb{C}$, we consider the normalized induced representation
$I(\lambda):={\rm Ind}^{G(\mathbb{R})}_{P(\mathbb{R})}\,|\nu|^{\lambda-\frac{n}{2}-1}.$
By the Peter-Weyl theorem, we have
${\rm Ind}^{K_{0}}_{K_{0}\cap P(\mathbb{R})}\mathbf{1}\simeq C^{\infty}((S^{2}\times S^{n})/\{\pm 1\})\simeq\bigoplus_{a,b\geq 0\atop a+b:\,{\rm even}}\mathcal{H}^{a}(\mathbb{C}^{3})\boxtimes\mathcal{H}^{b}(\mathbb{C}^{n+1})$
where $S^{k}$ stands for the $k$-dimensional sphere and $\mathcal{H}^{i}(\mathbb{C}^{j})$ stands for the harmonic polynomials on $\mathbb{C}^{j}$ of degree $i$. Using this, we have
(3.1) $I(\lambda)|_{K}=\bigoplus_{a,b\geq 0\atop a+b:\,{\rm even}}{\rm Ind}^{K}_{K_{0}}(\mathcal{H}^{a}(\mathbb{C}^{3})\boxtimes\mathcal{H}^{b}(\mathbb{C}^{n+1})).$
In particular, if $b=0$ (hence $a$ is even, say $a=2l$), then the action of $K$ factors through $O(3)$ and ${\rm Ind}^{K}_{K_{0}}(\mathcal{H}^{2l}(\mathbb{C}^{3})\boxtimes\mathbf{1})=\mathcal{H}^{2l}(\mathbb{C}^{3})\boxtimes\mathbf{1}$ appears in $I(\lambda)|_{K}$ with multiplicity one.
Definition 3.1.
Let $l\geq 0$ be an integer. Let $\mathbb{V}_{l}={\rm Sym}^{2l}(\mathbb{C}^{2})$ be the $(2l+1)$-dimensional representation of $K$ that factors through ${\rm O}(3)$ and we regard it with $\mathcal{H}^{2l}(\mathbb{C}^{3})\boxtimes\mathbf{1}$ as a $K$-type of $I(l+1)$. We fix a basis $\{[x]^{l+v}[y]^{l-v}\}_{v=-l}^{l}$ of $\mathbb{V}_{l}$ where $x^{k}]=\frac{x^{k}}{k!}$ and $y^{k}]=\frac{y^{k}}{k!}$ for $k\geq 0$. Modular forms on $G$ of weight $l$ are $\mathbb{V}_{l}$-valued automorphic functions $\phi$ on $G(\mathbb{A})$ that satisfy:
@@
(3) As a $(\mathfrak{g},K)$-module, $\phi$ generates an irreducible constituent of $I(l+1)$.
Cusp forms on $G$ of weight $l$ are similarly defined as in [6].
Proposition 3.2.
Assume $l>\frac{n}{2}$ is an even integer. Let $F$ be a non-zero modular form on $G$ of weight $l$ and $\Pi_{\infty}$ be the corresponding irreducible admissible representation of $G(\mathbb{R})$. Then, as a $(\mathfrak{g},K)$-module, $\Pi_{\infty}$ is isomorphic to $\Pi^{3,n+1}_{l,0}$ in the notation of [26, Proposition 2.1].
Proof.
By definition, $\Pi_{\infty}$ is a constituent of $I(l+1)$ with a $K$-type $\mathcal{H}^{2l}(\mathbb{C}^{3})\boxtimes\mathbf{1}$. Since $l>\frac{n}{2}$ and $l$ is even, by [26, Proposition 2.1], there is a unique irreducible submodule $\Pi^{3,n+1}_{l,0}$ of $I(l+1)$ with the minimal $K$-type $\mathcal{H}^{2l}(\mathbb{C}^{3})\boxtimes\mathbf{1}$. As observed, such a $K$-type is multiplicity free in $I(l+1)$. Thus, $\Pi_{\infty}$ and $\Pi^{3,n+1}_{l,0}$ share the same $K$-type. The irreducibility shows $\Pi_{\infty}\simeq\Pi^{3,n+1}_{l,0}$. ∎
@@
We embed $V^{\prime\prime}$ into $V^{\prime}$ by $\xi\mapsto(0,\xi,0)$ and write any element of $V^{\prime}$ as $\eta=(a,\xi,S)$ for $a,S\in\mathbf{G}_{a}$ and $\xi V^{\prime\prime}$.
Proposition 3.4.
Let $\phi$ be a modular form of weight $l\geq 1$ and be of level one, hence it is fixed by G(^)G($\widehat{\mathbb{Z}}$). If it has the Fourier expansion of the form
$\phi(g)=\sum_{\eta=(a,\xi,S)\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0,\ aS\neq 0}a_{\phi}(\eta)(g_{f})\mathcal{W}_{2\pi\eta}(g_{\infty}),$
then $\phi$ is a cusp form.
Proof.
By assumption, $G$ splits everywhere at finite places. Thus, we have $G(\mathbb{Q})=R(\mathbb{Q})G(\mathbb{Z})$ for any maximal standard $\mathbb{Q}$-parabolic subgroup $R$ of $G$. Therefore, we may check the cuspidality only for three maximal standard $\mathbb{Q}$-parabolic subgroups.
Since $G(\mathbb{A})=G(\mathbb{Q})(G(\widehat{\mathbb{Z}})\times G(\mathbb{R}))$, we may write
$\phi(g)=\sum_{\eta=(a,\xi,S)\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0,\ aS\neq 0}a_{\phi}(\eta)(1_{f})\mathcal{W}_{2\pi\eta}(g_{\infty}),\ g=\gamma k_{f}g_{\infty}.$
For each maximal standard $\mathbb{Q}$-parabolic subgroup $R$ with the unipotent radical $U_{R}$, by using the decomposition $G(\mathbb{A})=R(\mathbb{Q})(G(\widehat{\mathbb{Z}})\times G(\mathbb{R}))$, we have
$\int_{U_{R}(\mathbb{Q})\backslash U_{R}(\mathbb{A})}\phi(ug)du$ $=$ $\int_{U_{R}(\mathbb{Z})\backslash U_{R}(\mathbb{R})}\phi(u_{\infty}g_{\infty})du_{\infty}$
$=$ $\sum_{\eta=(a,\xi,S)\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0,\ aS\neq 0}a_{\phi}(\eta)(1_{f})\int_{U_{R}(\mathbb{Z})\backslash U_{R}(\mathbb{R})}\mathcal{W}_{2\pi\eta}(u_{\infty}g_{\infty})du_{\infty}.$
Thus, we may check the vanishing of $\int_{U_{R}(\mathbb{Z})\backslash U_{R}(\mathbb{R})}\mathcal{W}_{2\pi\eta}(u_{\infty}g_{\infty})du_{\infty}$ for each $R$. Since $q(\eta)\neq 0$, the constant term along $N$ (the unipotent radical of $R=P$) is vanishing. For other maximal parabolic subgroup $R$, it contains the center $Z$ of $U$ (here $U$ is the unipotent radical of $Q$). For each $z\in Z$, we see that $(\eta,z)_{V^{\prime}}=\frac{1}{2}az$. Thus, the constant term along $U_{R}$ is also vanishing since $a\neq 0$. ∎